ou have $2000 that you invested in an account t today. Interest is compounded monthly and you will withdraw $4000 in exactly 10 years. What is the effective interest rate?
Answer #1

This question requires application of basic time value of money function, according to which FV = PV * (1 + r) where FV is F

n = 10 * 12 = 120 months

$4000 = $2000 * (1 + r)120

2 = (1 + r)120

1.005793 = 1 + r

r = 0.5793%

This is monthly rate

Effective annual rate = (1 + 0.5793%)12 - 1= 1.0718 - 1 = 0.0718 = 7.18% --> Answer